# Can the 3 neutrino masses really be found using SN 1987A data?

### v5

@@ IV.12 Multiple sterile neutrinos in oscillation data?
Anatolii Serebrov and his Neutrino 4 team, have argued on cosmological grounds for there being two heavier sterile neutrinos in addition to the one found at $m_{1}^{2}=2.70^{2}eV^{2}/c^{4},$ which of course was also earlier inferred from the SN 1987A data analysis. Assuming a second sterile neutrino of mass $m_{2}^{2}=22.4eV^{2}/c^{2},$ as mandated by that data, the oscillation wavelength would be 69 times shorter than the $m_{1}$ sterile neutrinos. It is highly unlikely that such short wavelength oscillations are detectable in the data of Neutrino 4, but if they were one approach to find them would involve setting up a grid of values for the two masses $m_{1},m_{2}$ and in each case use the data to find the likelihood probability of the fit to the data for that pair of mass values. One could then see where on the plane that likelihood maximizes. If there is a peak right at a small region of the grid near $(m^{2}_{1},m_{2}^{2})=(2.70^{2},22.4^{2})eV^{2}/c^{2},$ that would constitute strong support for the idea that the $m^{2}$ masses of electron neutrinos consist of the values deduced from the SN 1987A data. You might think that the most promising experiments to look at would be the four data sets identified by Serebrov in ref. (Serebrov2024, ), which already have shown evidence of the $2.70eV/c^{2}$ mass at a cumulative $5.8\sigma$. Unfortunately, however, the very short wavelength of $22.4^{2}eV^{2}/c^{2}$ neutrinos almost certainly make these oscillations undetectable in existing short baseline reactor data. There may be over 20 projects in the works exploring the ultra-short baseline region ($L<10$ meters) in reactor experiments, (Serebrov2021, ) but no experiments known to the author could detect stable sterile neutrinos of such a high mass, with the exception of IceCube.
In fact, IceCube has sought evidence for $\nu_{\mu}$ and $\bar{\nu_{\mu}}$} disappearance oscillations in seven years of data. (Abbasi2024, ) Even though they reported a negative result their best fit did show a hint of a value consistent with the $m_{1}=2.70eV/c^{2}$ mass whuch was strongly preferred over the no oscillation solution by 10.7 to one. Furthermore, the IceCube energy range (up to $E=100TeV$) would make them sensitive to sterile mass values as high as $m_{2}^{2}=22.4^{2}eV^{2}/c^{4}.$ Thus, it might make sense for them to rexamine their data for evidence of a double oscillation for those two specific masses. Making such a search even more promissing, IceCube has already begun to explore the region of $\Delta m^{2}>10eV^{2}/c^{4},$ and found a preference for $3+1$ oscillations. (Biennow2018, ) A search using the method outlined for the Neutrino-4 data might prove particularly promising. Although the SN 1987A data analysis specified three neutrino masses it seems highly unlikely that the $m_{3}^{2}=-200,000eV^{2}/c^{4}$ neutrino would leave any detectable signal in IceCube or any other oscillation data (unlike KATRIN where it should be detectable).
@@ V Summary and the ultimate test
This paper has made six claims that many readers may understandably consider very difficult to believe – see Table 1. Some of the claims will likely not be accepted as true without the strongest corroborating evidence. After KATRIN does the analysis suggested in section IV.10, and it were to find evidence for the three active neutrino masses deduced from the fit to SN 1987 data, that might be enough. This might also be the case if oscillation experiments were to find such evidence doing the analysis mentioned in the previous section.
@@ 1. The LSD burst was due to SN 1987A but no 2nd bang
2. Upper limits on the 3 neutrino $m^{2}_{\nu,j}$ can be avoided
3. 3 $m_{j}^{2}$’s really can be found from 30 SN 1987A neutrinos
@@ 4. One of the 3 $m^{2}_{j}$ from SN 1987A data is a $m^{2}<0$ tachyon
5. The three $m^{2}=2.7^{2}eV^{2},22.4^{2}eV^{2}$ and $2\times 10^{5}eV^{2}$
6. KATRIN, IceCube or other data likely can confirm #5
Table 1: six claims made in this paper.
However, the most convincing test would be a new galactic supernova. Given the much larger neutrino detectors today compared to those in 1987, a new galactic supernova might yield thousands of neutrinos reaching the detectors. This 88 year-old does not have a great chance of seeing it, but let’s be an optimist. When supernova SN 2026(!) occurs at a distance that is a percentage p of the distance of SN 1987A, and we were to find an LSD-like near-monochromatic burst of a thousand neutrinos occurring earlier than the main neutrino burst by a time interval of $484\times p$ minutes, there would be little doubt that tachyons exist, as well as the other five claims. Even if the next galactic supernova does not occur in 2026, and I don’t get to see it, it is comforting to realize the evidence, in the form of thousands of neutrinos are now on their way to Earthly detectors.
These neutrinos will definitely either confirm or reject the six claims made. Moreover, given the known frequency of supernovae, we know that those neutrinos are now more than $99.99\%$ of their way here from some star in our galaxy. No one knows which star it will be, but Betelgeuse, a massive red supergiant in Orion, is expected to end its life in a spectacular supernova explosion soon. Soon in astronomical terms means potentially within the next 100,000 years (some say under 300 years (Saio2023, )). Were Betelgeuse to become a supernova, it would become incredibly bright, potentially as bright as the full moon. Go Betelgeuse!

### v6

@@ IV.12 Multiple sterile neutrinos in oscillation data?
The need for doing searches for multiple sterile neutrinos in a single search has already been noted in section III.3. How would one go about seeking evidence for both $m_{1}^{2}=2.70^{2}eV^{2}/c^{4},$ and $m_{2}=21.6eV^{2}/c^{4}$ in oscillation data, bearing in mind that the oscillation wavelength for $m_{2}$ would be $21.6^{2}/2.7^{2}=64$ times shorter than the $m_{1}$ sterile neutrinos? It is unlikely that such short wavelength oscillations are detectable in the data of Neutrino 4, but if they were, one approach to find them would involve setting up a grid of values for the two masses $m_{1},m_{2}$ and in each case use the data to find the likelihood probability of the fit to the data for that pair of mass values. One could then see if there is a peak right at a small region of the grid near $(m^{2}_{1},m_{2}^{2})=(2.70^{2},21.6^{2})eV^{2}/c^{2}.$ A more likely possibility would be the IceCube experiment. IceCube has sought evidence for $\nu_{\mu}$ and $\bar{\nu_{\mu}}$} disappearance oscillations in 10.7 years of data. (Abbasi2025, ) Even though IceCube reported a negative result their best fit they did show a hint of a value consistent with the $m_{1}=2.70eV/c^{2}$ mass which was strongly preferred over the no oscillation solution by 40 to one $p=2.5\%.$
@@ V Summary and the ultimate test
This paper has made five claims that many readers may understandably consider very difficult to believe – see Table 2. Some of the claims will likely not be accepted as true without the strongest corroborating evidence. After KATRIN does the analysis suggested in section IV.10, and it were to find evidence for the three active neutrino masses deduced from the fit to SN 1987 data, that might convince many skeptics.
@@ 1. The LSD burst was due to SN 1987A but no 2nd bang
2. 3 $m_{j}^{2}$’s really can be found from 37 SN 1987A neutrinos
@@ 3. One of the 3 $m^{2}_{j}$ from SN 1987A data is a $m^{2}<0$ tachyon
4. The three $m^{2}=2.7^{2}eV^{2},21.6eV^{2}$ and $1.92\times 10^{5}eV^{2}$
5. KATRIN and other data likely can confirm the above
Table 2: Five claims made in this paper.
However, the most convincing test would be a new galactic supernova. Given the much larger neutrino detectors today compared to those in 1987, a new galactic supernova might yield thousands of neutrinos reaching the detectors. This 88 year-old does not have a great chance of seeing it, but let’s be an optimist. When supernova SN 2026(!) occurs at a distance that is a percentage p of the distance of SN 1987A, and we were to find an LSD-like near-monochromatic burst of a thousand neutrinos occurring earlier than the main neutrino burst by a time interval of $484\times p$ minutes, there would be little doubt that tachyons exist, as well as the other four claims. Even if the next galactic supernova does not occur in 2026, and I don’t get to see it, it is comforting to realize the evidence, in the form of thousands of neutrinos are now on their way to Earthly detectors.
These neutrinos will definitely either confirm or reject the five claims made. Moreover, given the known frequency of supernovae, we know that those neutrinos are now more than $99.99\%$ of their way here from some star in our galaxy. No one knows which star it will be, but Betelgeuse, a massive red supergiant in Orion, is expected to end its life in a spectacular supernova explosion soon. Soon in astronomical terms means potentially within the next 100,000 years (some say under 300 years (Saio2023, )). Were Betelgeuse to become a supernova, it would become incredibly bright, potentially as bright as the full moon. Go Betelgeuse!